# From Text to [Uncaptioned image]: How PEFT-Driven Personality Manipulation Unleashes the Emoji Potential in LLMs

### v1

@@ Abstract
As the demand for human-like interactions with LLMs continues to grow, so does the interest in manipulating their personality traits, which has emerged as a key area of research. Methods like prompt-based In-Context Knowledge Editing (IKE) and gradient-based Model Editor Networks (MEND) have been explored but show irregularity and variability. IKE depends on the prompt, leading to variability and sensitivity, while MEND yields inconsistent and gibberish outputs. To address this, we employed Opinion QA Based Parameter-Efficient Fine-Tuning (PEFT), specifically Quantized Low-Rank Adaptation (QLORA), to manipulate the Big Five personality traits: Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism. After PEFT, models such as Mistral-7B-Instruct and Llama-2-7B-chat began generating emojis, despite their absence in the PEFT data. For instance, Llama-2-7B-chat generated emojis in 99.5% of extraversion-related test instances, while Mistral-8B-Instruct did so in 92.5% of openness-related test instances. Explainability analysis indicated that the LLMs used emojis intentionally to express these traits. This paper provides a number of novel contributions. First, introducing an Opinion QA dataset for PEFT-driven personality manipulation; second, developing metric models to benchmark LLM personality traits; third, demonstrating PEFT’s superiority over IKE in personality manipulation; and finally, analyzing and validating emoji usage through explainability methods such as mechanistic interpretability and in-context learning explainability methods. [^1] The code, models, and results will be made publicly available upon acceptance.
@@ 1 Introduction and Related Work
This paper addresses the challenges of personality manipulation in LLMs by introducing a novel approach grounded in the Big Five personality model. We present a new Opinion QA dataset and methodologies for systematically adjusting personality traits in LLMs. Utilising Quantized Low-Rank Adaptation (QLoRA), a method within Parameter-Efficient Fine-Tuning (PEFT) (dettmers2023qloraefficientfinetuningquantized), we demonstrate that LLMs can achieve more consistent and enduring personality expressions. This approach enhances model adaptability in interactions and reveals new behaviors, such as spontaneous emoji generation in Mistral-7B-Instruct and Llama-7B-chat, absent in the original models, following PEFT-based adjustments. Through mechanistic interpretability (Bereska and Gavves (2024)), we confirmed that pre-training data likely influenced emoji usage, while ICL (Brown (2020)) explainability verified that the emojis were not random artifacts. Our findings suggest this phenomenon represents a novel mode of expression linked to specific personality traits (Figure 1), introducing a new dimension of LLM communication that integrates verbal and visual elements.
@@ 2 Methodology
This paper explores manipulation in autoregressive transformer models using Parameter-Efficient Fine-Tuning (PEFT) and In-Context Knowledge Editing (IKE), focusing on Llama-2-7B-Chat (Touvron et al. (2023)), Llama-3-8B-Instruct (MetaAI (2023)), and Mistral-7B-Instruct (Jiang et al. (2023a).
@@
This work expands on the dataset from Mao et al. (2023), which consisted of opinion-based QA on specific topics, by better aligning with the topics and more accurately capturing personality traits. This study adds Openness and Conscientiousness to the original traits of Extraversion, Agreeableness, and Neuroticism. While Mao et al. (2023) excluded these dimensions, considering them similar to Agreeableness in generating viewpoints, we argue their inclusion is vital for a comprehensive analysis and understanding of trait influence on opinions. The dataset contains $5000$ instances, split $80:20$ for training ($4000$ instances, $800$ per trait) and testing ($1000$ instances, $200$ per trait). A GPT-3.5-based model generated opinion texts using structured prompts to elicit specific traits, enabling a nuanced analysis of personality expression. The text was manually verified to ensure alignment with intended traits. A multi-class personality classifier based on RoBERTa (Liu et al. (2019)), fine-tuned on the enhanced dataset, achieved $91.9\%$ accuracy on the test set. Classifier validation involved human verification of textual feature importance using SHAP (Lundberg and Lee (2017)) and LIME (Ribeiro et al. (2016)) to ensure predictions aligned with human understanding. Results were consistent, except for extraversion, where SHAP analysis was less representative due to the trait’s complexity. Further details about the dataset, classifier, and validation results are in A.1, A.2, and A.3.
@@
The In-Context Knowledge Editing (IKE) method (Zheng et al. (2023)) was used as a baseline to manipulate embedded knowledge in LLMs, serving as a comparative foundation for evaluating prompt-based versus fine-tuning techniques in personality manipulation. The same prompt from Mao et al. (2023), provided in A.6, ensured consistency in comparing IKE with PEFT. We excluded methods like MEND due to inconsistent and gibberish outputs. Parameter-Efficient Fine-Tuning (PEFT), specifically Quantized Low-Rank Adaptation (QLORA), was selected to reduce computational cost while maintaining performance. The process began by preparing the Personality dataset, formatting an additional column as <s>[INST] Question [/INST] Answer </s>. Models were loaded from Hugging Face using 4-bit quantization via BitsAndBytes (Dettmers et al. (2023)) for efficient processing and storage. The temperature was set to 1. The hyperparameters, chosen for their balance of efficiency and performance (Houlsby et al. (2019); Dettmers et al. (2024)), are detailed in A.4. After training, fine-tuned LoRa weights were merged with the original model for evaluation.
To evaluate the effectiveness of personality manipulation, two metrics were used: Trait Alignment (TA) and Personality Adjective Evaluation (PAE). TA is computed using the metric classifier by comparing predicted labels $\hat{y}_{i}$ with true labels $y_{i}$, which correspond to the original target personality traits from the dataset. For a dataset with $N$ instances, TA is given by $\text{TA}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{1}(\hat{y}_{i}=y_{i})$, where $\mathbb{1}(\hat{y}_{i}=y_{i})$ is 1 if the prediction matches the true label and 0 otherwise, providing an average alignment between predictions and the intended personality traits. PAE, inspired by Mao et al. (2023), uses Chain of Thought (CoT) prompting (Wei et al. (2022)), where a larger LLM (here GPT-4 (Brown (2020))) scores generated text on a 1-5 scale based on its alignment with the target trait. The PAE score is calculated as the difference between the score of generated text and original text from the dataset, with higher scores indicating better alignment. The final PAE is the mean of these score differences across all instances, $\text{PAE}=\frac{1}{N}\sum_{i=1}^{N}s_{i}$, where $s_{i}$ is the score difference for each instance $i$. Details of the prompt are provided in A.5.
@@ 3 Result and Discussion
The performance of PEFT and IKE is presented in Figure 2 and detailed in Table 11 in A.9. PEFT consistently outperforms IKE in TA across most personality traits and models, indicating more reliable manipulation. Although IKE occasionally exceeds PEFT in PAE for certain traits by capturing nuanced text alignment, PEFT provides deeper and more stable personality manipulation and better scalability when applied across multiple traits. During inference, fine-tuned Llama-2-7B-Chat and Mistral-7B-Instruct models generated emojis, despite no emojis being present in the PEFT training data. In contrast, the original models produced generic responses without emojis for the same inputs. This behavior is likely due to pre-training on diverse corpora containing emoji patterns, with PEFT amplifying these latent tendencies. To test this, we performed neuron activation analysis, a mechanistic interpretability method (Bereska and Gavves (2024)), finding that specific neurons in Mistral-7B-Instruct and Llama-2-7B-Chat were responsible for emoji generation, becoming more active during conversational prompts. This suggests pre-training on informal data influenced these behaviors. In contrast, Llama-3-8B-Instruct showed no such activation, implying that these tendencies were either not learned or suppressed during fine-tuning (see A.7). To verify that the emojis were intentional and not random artefacts, we also conducted an in-context learning (ICL) explainability analysis through prompting, confirming that emoji usage aligned with intended personality traits (further in A.8). The Emoji-to-Sentence Ratio (ESR) was used to measure the frequency of emojis in generated responses after PEFT, calculated as ESR=Sentences with emojisTotal $\text{ESR}=\frac{\text{Sentences with emojis}}{\text{Total sentences}}$. Table 1 shows the results from mechanistic interpretability and ICL analysis, highlighting frequently used emojis and their corresponding ESRs. Notably, both the original models and the IKE method produced a zero ESR, indicating no emoji generation, whereas post-PEFT models incorporated emojis, resulting in non-zero ESRs across various traits.
@@
Limitations and Future Work: This study assesses personality manipulation in LLMs using the Big Five model, revealing inconsistencies such as high TA in Neuroticism (0.975) and low TA in Agreeableness (0.065) in Llama-2-7B-Chat. The disparity may stem from Neuroticism’s emotional salience versus Agreeableness’s context dependence. Future work should enhance fine-tuning consistency and consider continuous personality measures to capture nuances missed by classifiers. Additionally, exploring alternative psychometric models like 16PF Cattell and Mead (2008) could provide further insights into personality manipulation.
@@ A.1 Personality Manipulation Dataset
For instance, as visible in the figure 4, the word cloud for extraversion showcases words that reflect the energetic, outgoing, and sociable nature of individuals with this personality trait. Prominent terms such as "absolutely," "always," "believe," "love," and "thrilling" emphasise the enthusiasm, positivity, and social engagement typical of extraverts. Similarly, the word cloud for neuroticism reveals words associated with emotional instability, anxiety, and self-consciousness. Key terms like "feel," "worry," "sometimes," "make," and "bit" are prominently displayed, reflecting the frequent emotional fluctuations and concerns typical of neurotic individuals. Words like "anxious," "overwhelming," and "struggle" further emphasise the challenges faced by those with high levels of neuroticism. Collectively, these word clouds provide a rich visual representation of the lexical diversity associated with each personality trait, offering valuable insights into the distinct characteristics and behavioral tendencies that define them.
@@ A.1.1 Text Analysis
In conclusion, LDA-based topic modeling reveals significant insights into how different themes resonate with various personality traits. This understanding is crucial for the manipulation of personalities in language models, allowing for the creation of content that is tailored to engage specific personality profiles effectively. By leveraging these insights, content can be crafted to not only align with individual preferences but also to influence and shape personality-driven responses and behaviors. This approach enhances personalised communication strategies and fosters deeper, more meaningful connections with diverse audiences.
@@ A.7 Neuron Activation Analysis
To support our hypothesis that PEFT amplifies subtle, pre-existing emoji patterns learned from pre-training on diverse training corpora, we conducted Neuron Activation Analysis. This analysis used conversational and informal prompts to trigger potential emoji-related activations in deepest transformer layer just before the output layer, which generates the final token predictions, providing further insights into the underlying behavior.
@@
In Llama-3-8B-Instruct, the prompts fail to activate neurons responsible for emoji generation, suggesting that these behaviors were either not learned during pre-training or suppressed during fine-tuning. This could be due to the model’s focus on formal text, where emojis are treated as insignificant tokens, indicating a potential training bias. Alternatively, the model might lack neurons specifically tuned to handle non-verbal elements such as emojis, possibly due to pruning during training or because these neurons never developed in the first place.
@@
As seen in figure 14, for Llama-2-7B-chat there may be certain neurons that are highly specialised in generating or recognising emojis. This is indicated by the sharp peak in neuron activation for the emoji "[Uncaptioned image]". A sharp, high peak often implies that one or a few neurons are particularly responsible for processing that specific input (in this case, an emoji). This suggests that the model has dedicated neurons for handling emoji-related behavior, and those neurons activate strongly when encountering emojis in conversational prompts.
@@
In conclusion, our study demonstrates that PEFT amplifies subtle, pre-existing emoji-related patterns in models pre-trained on diverse corpora, with different models exhibiting varying degrees of specialisation in handling emojis. Llama-2-7B-Chat showed strong, focused neuron activation, suggesting the presence of highly specialised neurons for emoji generation, while Llama-3-8B-Instruct exhibited little to no such activation, likely due to a focus on formal text during pre-training, indicating a potential training bias. Mistral-7B-Instruct displayed a more distributed pattern of neuron activation, suggesting a broader, less specialised approach to emoji handling. These findings highlight how model architecture and training data influence the expression of informal behaviors like emoji usage, with PEFT further amplifying these tendencies where they exist.
A.8 Manipulation Validation

### v2

# From Text to Emoji: How PEFT-Driven Personality Manipulation Unleashes the Emoji Potential in LLMs

@@ Abstract
As the demand for human-like interactions with LLMs continues to grow, so does the interest in manipulating their personality traits, which has emerged as a key area of research. Methods like prompt-based In-Context Knowledge Editing (IKE) and gradient-based Model Editor Networks (MEND) have been explored but show irregularity and variability. IKE depends on the prompt, leading to variability and sensitivity, while MEND yields inconsistent and gibberish outputs. To address this, we employed Opinion QA Based Parameter-Efficient Fine-Tuning (PEFT), specifically Quantized Low-Rank Adaptation (QLoRA), to manipulate the Big Five personality traits: Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism. After PEFT, models such as Mistral-7B-Instruct and Llama-2-7B-chat began generating emojis, despite their absence in the PEFT data. For instance, Llama-2-7B-chat generated emojis in 99.5% of extraversion-related test instances, while Mistral-7B-Instruct did so in 92.5% of openness-related test instances. Explainability analysis indicated that the LLMs used emojis intentionally to express these traits. This paper provides a number of novel contributions. First, introducing an Opinion QA dataset for PEFT-driven personality manipulation; second, developing metric models to benchmark LLM personality traits; third, demonstrating PEFT’s superiority over IKE in personality manipulation; and finally, analysing and validating emoji usage through explainability methods such as mechanistic interpretability and in-context learning explainability methods.
@@ 1 Introduction and Related Work
This paper addresses the challenges of personality manipulation in LLMs by introducing a novel approach grounded in the Big Five personality model. We present a new Opinion QA dataset and methodologies for systematically adjusting personality traits in LLMs. Utilising Quantized Low-Rank Adaptation (QLoRA), a method within Parameter-Efficient Fine-Tuning (PEFT) (Dettmers et al. (2023a)), we demonstrate that LLMs can achieve more consistent and enduring personality expressions. This approach enhances model adaptability in interactions and reveals new behaviours, such as spontaneous emoji generation in Mistral-7B-Instruct and Llama-7B-chat, absent in the original models, following PEFT-based adjustments. Through mechanistic interpretability (Bereska and Gavves (2024)), we confirmed that pre-training data likely influenced emoji usage, while ICL (Brown (2020)) explainability verified that the emojis were not random artifacts. Our findings suggest this phenomenon represents a novel mode of expression linked to specific personality traits (Figure 1), introducing a new dimension of LLM communication that integrates verbal and visual elements which enhances user engagement, improves emotional expressiveness in digital assistants, and enables more personalized user experiences in areas such as mental health, education, and customer service (Votintseva et al. (2024)).
@@ 2 Methodology
This paper explores manipulation in autoregressive transformer models using Parameter-Efficient Fine-Tuning (PEFT) and In-Context Knowledge Editing (IKE), focusing on Llama-2-7B-Chat (Touvron et al. (2023)), Llama-3-8B-Instruct (MetaAI (2023)), and Mistral-7B-Instruct (Jiang et al. (2023a)).
@@
This work expands on the dataset from Mao et al. (2023), which consisted of opinion-based QA on specific topics, by better aligning with the topics and more accurately capturing personality traits. This study adds Openness and Conscientiousness to the original traits of Extraversion, Agreeableness, and Neuroticism. While Mao et al. (2023) excluded these dimensions, considering them similar to Agreeableness in generating viewpoints, we argue their inclusion is vital for a comprehensive analysis and understanding of trait influence on opinions. The dataset contains $5000$ instances, split $80:20$ for training ($4000$ instances, $800$ per trait) and testing ($1000$ instances, $200$ per trait).  [^1] Dataset: https://huggingface.co/datasets/holistic-ai/personality_manipulation. A GPT-3.5-based model generated opinion texts using structured prompts to elicit specific traits, enabling a nuanced analysis of personality expression. The generated text was analysed using word clouds and text analysis to identify key linguistic patterns, thematic elements, and ensure lexical diversity associated with the Big Five personality traits, as detailed in Appendix A.1. The text was manually verified to ensure alignment with intended traits. A multi-class personality classifier based on RoBERTa (Liu et al. (2019)), fine-tuned on the enhanced dataset, achieved $91.9\%$ accuracy on the test set.  [^2] Classifier: https://huggingface.co/holistic-ai/personality_classifier. Classifier validation involved human verification of textual feature importance using SHAP (Lundberg and Lee (2017)) and LIME (Ribeiro et al. (2016)) to ensure predictions aligned with human understanding. For instance, in the sentence, "I believe the First Indochina War had its consequences, paving the way for the withdrawal of French colonial forces. However, there were many factors at play, and it’s important to acknowledge the contributions of everyone involved." certain words play a pivotal role in predicting the Agreeableness trait as observed in Figures 2 and 3. Terms such as "contributions", "acknowledge", and "believe" have a strong positive contribution to the prediction of Agreeableness, as they suggest inclusiveness, recognition, and a conciliatory tone, which are characteristic of Agreeableness. On the other hand, words like "consequences" contribute negatively to the Agreeableness prediction because it often connotes conflict, repercussions, or negative outcomes, whereas Agreeableness is characterized by cooperation, empathy, and a focus on harmony and positive social interactions (Liu and Sun (2020)).
Figure 2: SHAP visualisation for Agreeableness
Figure 3: LIME visualisation for Agreeableness
Thus, as seen above and in other examples inA.3, the results were consistent, except for extraversion, where SHAP analysis was less representative due to the trait’s complexity. Through above analysis, we can conclude that the personality classifier has successfully captured the expected patterns from perspective of feature attribution. Further details about the classifier are in A.2.
@@
The In-Context Knowledge Editing (IKE) method (Zheng et al. (2023)) was used as a baseline to manipulate embedded knowledge in LLMs, serving as a comparative foundation for evaluating prompt-based versus fine-tuning techniques in personality manipulation. The same prompt from Mao et al. (2023), provided in A.6, ensured consistency in comparing IKE with PEFT. We excluded methods like MEND due to inconsistent and gibberish outputs. Parameter-Efficient Fine-Tuning (PEFT), specifically Quantized Low-Rank Adaptation (QLoRA), was selected to reduce computational cost while maintaining performance. The process began by preparing the Personality dataset, formatting an additional column as <s>[INST] Question [/INST] Answer </s>. Models were loaded from Hugging Face using 4-bit quantization via BitsAndBytes (Dettmers et al. (2023b)) for efficient processing and storage. The temperature was set to 1. The hyperparameters, chosen for their balance of efficiency and performance (Houlsby et al. (2019); Dettmers et al. (2024)), are detailed in A.4. After training, fine-tuned LoRa weights were merged with the original model for evaluation.
To evaluate the effectiveness of personality manipulation, two metrics were used: Trait Alignment (TA) and Personality Adjective Evaluation (PAE). TA is computed using the metric classifier by comparing predicted labels $\hat{y}_{i}$ with true labels $y_{i}$, which correspond to the original target personality traits from the dataset. For a dataset with $N$ instances, TA is given by $\text{TA}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{1}(\hat{y}_{i}=y_{i})$, where $\mathbb{1}(\hat{y}_{i}=y_{i})$ is 1 if the prediction matches the true label and 0 otherwise, providing an average alignment between predictions and the intended personality traits. PAE, inspired by Mao et al. (2023), uses Chain of Thought (CoT) prompting (Wei et al. (2022)), where a larger LLM (here GPT-4 (Brown (2020))) scores generated text on a 1-5 scale based on its alignment with the target trait. The PAE score is calculated as the difference between the score of generated text and original text from the dataset, with higher difference indicating that the generated text after PEFT more effectively captures the intended personality traits compared to the original text. The final PAE is the mean of these score differences across all instances, $\text{PAE}=\frac{1}{N}\sum_{i=1}^{N}s_{i}$, where $s_{i}$ is the score difference for each instance $i$. Details of the prompt are provided in A.5.
@@ 3 Result and Discussion
The performance of PEFT and IKE is presented in Figure 4 and detailed in Table 11 in A.9. PEFT consistently outperforms IKE in TA across most personality traits and models, indicating more reliable manipulation. Although IKE occasionally exceeds PEFT in PAE for certain traits by capturing nuanced text alignment, PEFT provides deeper and more stable personality manipulation and better scalability when applied across multiple traits. We additionally employed prompting and manual verification to validate effective manipulation across all traits (in A.8). During inference, fine-tuned Llama-2-7B-Chat and Mistral-7B-Instruct models generated emojis, despite no emojis being present in the PEFT training data. In contrast, the original models produced generic responses without emojis for the same inputs. This behaviour is likely due to pre-training on diverse corpora containing emoji patterns, with PEFT amplifying these latent tendencies. To test this, we performed neuron activation analysis, a mechanistic interpretability method (Bereska and Gavves (2024)), finding that specific neurons in Mistral-7B-Instruct and Llama-2-7B-Chat were responsible for emoji generation, becoming more active during conversational prompts. This suggests pre-training on informal data influenced these behaviours. In contrast, Llama-3-8B-Instruct showed no such activation, implying that these tendencies were either not learned or suppressed during fine-tuning when first developed (see A.7). To verify that the emojis were intentional and not random artefacts, we also conducted an in-context learning (ICL) explainability analysis through prompting, confirming that emoji usage aligned with intended personality traits. This analysis followed the same method as manipulation result validation (in A.8) but focused specifically on emojis. The Emoji-to-Sentence Ratio (ESR) was used to measure the frequency of emojis in generated responses after PEFT, calculated as ESR=Sentences with emojisTotal $\text{ESR}=\frac{\text{Sentences with emojis}}{\text{Total sentences}}$. Table 1 shows the results from mechanistic interpretability and ICL analysis, highlighting frequently used emojis and their corresponding ESRs. Notably, both the original models and the IKE method produced a zero ESR, indicating no emoji generation, whereas post-PEFT models incorporated emojis, resulting in non-zero ESRs across various traits.
@@
4 Limitations and Future Work
This study assesses personality manipulation in LLMs using the Big Five model, revealing inconsistencies such as high TA in Neuroticism (0.975) and low TA in Agreeableness (0.065) in Llama-2-7B-Chat. The disparity may stem from Neuroticism’s emotional salience versus Agreeableness’s context dependence. Future work should enhance fine-tuning consistency and consider continuous personality measures to capture nuances missed by classifiers. Additionally, exploring alternative psychometric models like 16PF (Cattell and Mead (2008)) could provide further insights into personality manipulation. Future work could also explore how model size correlates with manipulation effectiveness and whether there is a threshold where scaling yields diminishing returns.
@@ A.1 Personality Manipulation Dataset
For instance, as visible in the figure 6, the word cloud for extraversion showcases words that reflect the energetic, outgoing, and sociable nature of individuals with this personality trait. Prominent terms such as "absolutely," "always," "believe," "love," and "thrilling" emphasise the enthusiasm, positivity, and social engagement typical of extraverts. Similarly, the word cloud for neuroticism reveals words associated with emotional instability, anxiety, and self-consciousness. Key terms like "feel," "worry," "sometimes," "make," and "bit" are prominently displayed, reflecting the frequent emotional fluctuations and concerns typical of neurotic individuals. Words like "anxious," "overwhelming," and "struggle" further emphasise the challenges faced by those with high levels of neuroticism. Collectively, these word clouds provide a rich visual representation of the lexical diversity associated with each personality trait, offering valuable insights into the distinct characteristics and behavioural tendencies that define them.
@@ A.1.1 Text Analysis
In conclusion, LDA-based topic modeling reveals significant insights into how different themes resonate with various personality traits. This understanding is crucial for the manipulation of personalities in language models, allowing for the creation of content that is tailored to engage specific personality profiles effectively. By leveraging these insights, content can be crafted to not only align with individual preferences but also to influence and shape personality-driven responses and behaviours. This approach enhances personalised communication strategies and fosters deeper, more meaningful connections with diverse audiences.
@@ A.7 Neuron Activation Analysis
To support our hypothesis that PEFT amplifies subtle, pre-existing emoji patterns learned from pre-training on diverse training corpora, we conducted Neuron Activation Analysis. This analysis used conversational and informal prompts to trigger potential emoji-related activations in deepest transformer layer just before the output layer, which generates the final token predictions, providing further insights into the underlying behaviour (Deng et al. [2024]).
@@
In Llama-3-8B-Instruct, the prompts fail to activate neurons responsible for emoji generation, suggesting that these behaviours were either not learned during pre-training or suppressed during fine-tuning. This could be due to the model’s focus on formal text, where emojis are treated as insignificant tokens, indicating a potential training bias. Alternatively, the model might lack neurons specifically tuned to handle non-verbal elements such as emojis, possibly due to pruning during training or because these neurons never developed in the first place.
@@
As seen in figure 16, for Llama-2-7B-chat there may be certain neurons that are highly specialised in generating or recognising emojis. This is indicated by the sharp peak in neuron activation for the emoji "[Uncaptioned image]". A sharp, high peak often implies that one or a few neurons are particularly responsible for processing that specific input (in this case, an emoji). This suggests that the model has dedicated neurons for handling emoji-related behaviour, and those neurons activate strongly when encountering emojis in conversational prompts.
@@
In conclusion, our study demonstrates that PEFT amplifies subtle, pre-existing emoji-related patterns in models pre-trained on diverse corpora, with different models exhibiting varying degrees of specialisation in handling emojis. Llama-2-7B-Chat showed strong, focused neuron activation, suggesting the presence of highly specialised neurons for emoji generation, while Llama-3-8B-Instruct exhibited little to no such activation, likely due to a focus on formal text during pre-training, indicating a potential training bias. Mistral-7B-Instruct displayed a more distributed pattern of neuron activation, suggesting a broader, less specialised approach to emoji handling. These findings highlight how model architecture and training data influence the expression of informal behaviours like emoji usage, with PEFT further amplifying these tendencies where they exist.
A.8 Manipulation Validation and ICL Explainability